OpenStudy (anonymous):

If z is directly proportional to y^3 and z=1/9 when y=1/3, what is the value of y when z=24?

OpenStudy (anonymous):

z is directly proportional to y z=cy^3 where c is constant of proportionality when z=1/9 and y=1/3 c=3 when z=24, 24=3*y^3 y^3=8 y=2